Solve the equation in the complex number system:
65x^2 +1 =8x
I'm so lost on how to even begin this!

Answer by stanbon(75887)

Solve the equation in the complex number system:
65x^2 +1 =8x
-----
65x^2-8x+1 = 0
-------
Quadratic Formula:
x = [8 +- sqrt(64-4*65*1)]/130
--------
x = [8 +- sqrt(-196)]/130
---
x = [8 +- 14i]/130
---
x = 4/65 +- (7/65)i
========================
Cheers,
Stan H.
==============